### v1

# GFlowGR: Fine-tuning Generative Recommendation Frameworks with Generative Flow Networks

Abstract.
Generative recommendations (GR), which usually include item tokenizers and generative Large Language Models (LLMs), have demonstrated remarkable success across a wide range of scenarios. The majority of existing research efforts primarily concentrate on developing powerful item tokenizers or advancing LLM decoding strategies to attain superior performance. However, the critical fine-tuning step in GR frameworks, which is essential for adapting LLMs to recommendation data, remains largely unexplored. Current approaches predominantly rely on either the next-token prediction loss of supervised fine-tuning (SFT) or recommendation-specific direct preference optimization (DPO) strategies. Both methods ignore the exploration of possible positive unobserved samples, which is commonly referred to as the exposure bias problem. To mitigate this problem, this paper treats the GR as a multi-step generation task and constructs a GFlowNets-based fine-tuning framework (GFlowGR). The proposed framework integrates collaborative knowledge from traditional recommender systems to create an adaptive trajectory sampler and a comprehensive reward model. Leveraging the diverse generation property of GFlowNets, along with sampling and heuristic weighting techniques, GFlowGR emerges as a promising approach to mitigate the exposure bias problem. Extensive empirical results on two real-world datasets and with two different GR backbones highlight the effectiveness and robustness of GFlowGR.
Recommender Systems; Generative Recommendations
$\dagger$ Work was conducted during the internship of Yejing Wang at Alibaba.
* Xiangyu Zhao is corresponding author.
copyright: acmcopyrightjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06
1. Introduction
Large Language Models (LLMs) possess exceptional capabilities in understanding, reasoning, and generation. As a result, they are increasingly utilized across various domains, including healthcare, education, and science (Naveed et al., 2023), as well as for information retrieval and recommendations in numerous online services, such as E-commerce (Li et al., 2024a; Zhai, 2024). Notably, the impressive generative capabilities of LLMs have spurred the development of generative recommendation (GR) frameworks (Rajput et al., 2023; Huang et al., 2024; Hua et al., 2023). As visualized in Figure 1 (a), these frameworks typically consist of three key components (Li et al., 2024b): item tokenizers, user prompts, and generative LLMs. Meanwhile, the training paradigms usually consist of two stages: optimizing the item tokenizer to represent items with several tokens and fine-tuning the LLMs for recommendation objectives like next-item prediction.
Existing research on GR has primarily concentrated on enhancing the item tokenizer (Zheng et al., 2025; Liu et al., 2024c; Lin et al., 2025b; Wang et al., 2024a; Zhu et al., 2024) and the associated optimization processes, designing user prompts (Liu et al., 2024a; Wang et al., 2024b), and advancing generative LLMs (Lin et al., 2025c; Ding et al., 2024). However, there has been little focus on innovating the fine-tuning step of LLMs. While this step is crucial for GR frameworks: the semantics of recommendations can significantly differ from those of LLMs pretrained for general language purposes (Lin et al., 2025a; Wu et al., 2024). For instance, terms like “Comment” represent specific user behaviors within recommendation data, even though they are commonly used verbs in general language tasks. Moreover, the item identifier in GR may consist of out-of-vocabulary tokens for pretrained LLMs (Huang et al., 2024; Wang et al., 2024a; Rajput et al., 2023), which are unlikely to be generated by zero-shot LLMs.
Current LLM fine-tuning techniques applied for GR can be generally divided into two categories (Wu et al., 2025; Wang et al., [n. d.]): supervised fine-tuning (SFT) and reinforcement learning-based fine-tuning (RLFT). The former optimizes LLMs with the next token prediction loss (Brown et al., 2020; Radford et al., 2019; Hu et al., 2022) and handcrafted auxiliary alignment loss (Liu et al., 2024a; Zheng et al., 2024; Wang et al., 2024b), which focuses solely on positive samples within the dataset. On the one hand, SFT does not incorporate any negative samples for comparison with the positive ones, leaving a gap with the ranking objective of recommender systems (Chen et al., [n. d.]). On the other hand, SFT cannot explore potential positive samples to extend the collected data, which is always sparse (Gao et al., 2022). With the exploration ability, RLFT methods like proximal policy optimization (PPO) (Schulman et al., 2017), and direct preference optimization (DPO) (Rafailov et al., 2023) can improve the model capacity on the basis of SFT by achieving the maximum reward. For example, recent DPO-based methods facilitate LLM fine-tuning by integrating augmented negative samples for GR (Chen et al., [n. d.]; Bai et al., 2024; Liao et al., 2024; Gao et al., 2025a; Deng et al., 2025), enabling LLMs to rank candidate items. However, both solutions cannot match the diverse nature of recommender systems that users may prefer multiple items, while only one among them can be observed. SFT regards the unobserved interactions as negative, and RLFT methods predefine the largest reward to the observed samples. This reliance on the collected data that fails to fully reveal the real user preference can lead to biased user preferences, known as exposure bias (Chen et al., 2023). Therefore, it calls for an LLM fine-tuning solution that can explore potential positive samples and leverage augmented samples to achieve diverse results.
Recent advancements in Generative Flow Networks (GFlowNets) present a promising solution: by training LLMs to generate token sequences with probabilities proportional to a specific reward distribution (Bengio et al., 2023), GFlowNets can identify potential positive candidates and thus extend the recommendation diversity. This capability is supported by recent studies demonstrating the application of GFlowNets for multi-step reasoning tasks (Kang et al., 2025; Takase et al., 2024; Yu et al., 2024) and compositional generation (Jain et al., 2022). Generative recommendation tasks are inherently similar to these tasks; as illustrated in Figure 1 (b), recommending an item involves multiple generation steps of a single token.However, applying GFlowNets for GR still faces significant challenges: how to organize training trajectories (samples) and establish a reward model to assess the terminal state for GFlowNets. Specifically, the collected data consists solely of positive interaction logs from users, resulting in sparse training trajectories for GFlowNets. This scenario necessitates high-quality augmented trajectories for robust training. Furthermore, merely treating these augmented trajectories as negative samples and assigning them low rewards is inadequate for overcoming the exposure bias problem, as these augmented trajectories are not consistently negative.
To address these limitations, this paper constructs the GFlowNets-based fine-tuning paradigm for GR (GFlowGR). Drawing inspiration from prior insights on aligning LLMs with collaborative recommendation models (Hong et al., [n. d.]; Yang et al., 2024; Lin et al., 2025a), we incorporate collaborative scores when designing the trajectory sampling and reward model for GFlowGR. Specifically, we develop an adaptive trajectory sampler that generates easy-to-hard augmented trajectories, thereby implementing a curriculum-style training approach (Wang et al., 2021a) to consistently provide high-quality training data. Additionally, we construct a multi-faceted reward system that thoroughly considers augmentation signals, collaborative scores, and semantic similarities to effectively differentiate possibly preferred items from the sampled unobserved trajectories. Finally, to further mitigate the exposure bias problem, we assign confidence weights to each sample based on the agreement between augmentation signals and collaborative scores. We summarize the major contributions of this paper as follows:
  • We formulate GR as a multi-step generation process and develop a GFlowNets-based fine-tuning framework, GFlowGR. This framework helps mitigate the exposure bias problem by learning generation probabilities that are proportional to the reward.
  • We implement an adaptive trajectory sampler to select high-quality training samples for GFlowGR and establish a comprehensive reward model to consider various reward signals effectively.
  • Our extensive experimental results, obtained from two public benchmarks using two backbone GR models, demonstrate the effectiveness and versatility of GFlowGR.
Figure 1. Overview of GR and GFlowNets, with the length of item identifiers $L=3$.
2. Preliminary
Generative Recommendation (GR). In this paper, we define GR frameworks by referencing recent works (Rajput et al., 2023; Wang et al., 2024a). A typical GR workflow is illustrated in Figure 1 (a). The framework first constructs an item tokenizer using item features to decompose items $v$ into token sequences $\{t_{l}\}_{l<L}$, serving as item identifiers. The lower portion of the figure exemplifies that a skyblue shirt is encoded as the token sequence $<a\_1><b\_4><c\_2>$ based on its title and descriptions. To infer the user preference, GR frameworks assemble user prompts (LLM inputs) by integrating task descriptions, historical user interactions, and profile information (e.g., age, gender denoted as $<\underline{X}>$). These prompts, exemplified by tokenized interaction histories (blue tokens for jackets, shirts, pants) and profiles, are fed into LLMs to generate token sequences representing the most probable next item. As shown in the upper portion of Figure 1 (a), this process produces $<a\_2><b\_4><c\_10>$, decoding to a blue polo shirt as the recommended item.
To achieve the recommendation goal, the training paradigm of GR comprises two stages: optimizing the item tokenizer ($\mathcal{L}_{T}$) and the generative LLMs ($\mathcal{L}_{\text{GR}}$), respectively. For the tokenizer, one prevalent tokenization strategy is based on RQ-VAE solutions (Lee et al., 2022; Rajput et al., 2023), where $\mathcal{L}_{T}$ includes a codebook loss and a reconstruction loss. Studies enhancing the tokenizer often incorporate auxiliary losses into $\mathcal{L}_{T}$, such as diversity (Wang et al., 2024a) and collaborative-constrained losses (Lin et al., 2025b). $\mathcal{L}_{\text{GR}}$ typically adopts the next-token prediction loss (Deng et al., 2025; Rajput et al., 2023; Wang et al., 2024a; Liu et al., 2024a)::::
(1) $\mathcal{L}_{\text{GR}}(U,v)=-\sum_{l=1}^{L}\mathrm{log}\mathbb{P}_{\text{GR}}(t_{l}|U,\{t_{i}\}_{i<l})$
where $L$ denotes the length of item identifiers (e.g., $L=3$ for the visualized examples in this paper). We denote the probability that LLMs generate the next ground-truth token $t_{l}$ given the prefix $U,\{t_{i}\}_{i<l}$ as $\mathbb{P}\left(t_{l}\,\big{|}\,U,\{t_{i}\}_{i<l}\right)$. Here, $\{t_{i}\}_{i<l}$ represents the item identifier tokens preceding $t_{l}$.
Generative Flow Networks (GFLowNets). The core idea of Generative Flow Networks (GFlowNets) stems from the generation of compositional objects through a sequence of stochastic steps. We formally define the state transition sequence as $\tau=(s_{0}\to s_{1}\to\dots\to s_{L})$, where $s_{L}$ represents the terminal generation state and $\tau$ denotes a specific generation trajectory. The training objective of GFlowNets is to align the generation probability of any trajectory $\tau$ with the reward obtained at its terminal state, such that $\mathbb{P}(\tau)\propto\mathcal{R}(s_{L})$  (Bengio et al., 2021; Takase et al., 2024). To achieve this, GFlowNets introduce a flow estimator $\mathcal{F}(s_{l})$ to model the likelihood of traversing state $s_{l}$, transforming the proportionality objective into a flow-matching principle: the incoming flow to a state must equal the outgoing flow. This is formalized as:
(2) $\mathcal{F}(s_{l})\mathbb{P}_{F}(s_{l+1}|s_{l})=\mathcal{F}(s_{l+1})\mathbb{P}_{B}(s_{l}|s_{l+1})$
Here, $\mathbb{P}_{F}(s_{l+1}|s_{l})$ denotes the forward transition probability from state $s_{l}$ to $s_{l+1}$, while $\mathbb{P}_{B}(s_{l}|s_{l+1})$ represents the backward probability that state $s_{l+1}$ originated from $s_{l}$. This bidirectional flow balance ensures that trajectories with higher terminal rewards $\mathcal{R}(s_{L})$ are assigned greater generation probabilities, enabling efficient exploration of complex combinatorial spaces (Bengio et al., 2023).
To facilitate practical implementation, two optimization-friendly variants are derived from the flow-matching principle (Malkin et al., 2022; Bengio et al., 2021): detailed balance (DB) and trajectory balance (TB). The former focuses on stepwise flow equilibrium, while the latter enforces balance at the full-trajectory level. Their respective loss functions are defined:
(3) $\mathcal{L}_{\text{DB}}(\tau)=\sum_{l=0}^{L-1}\Big{(}\mathrm{log}\frac{\mathcal{F}(s_{l})\mathbb{P}_{F}(s_{l+1}|s_{l})}{\mathcal{F}(s_{l+1})\mathbb{P}_{B}(s_{l}|s_{l+1})}\Big{)}^{2}$
(4) $\mathcal{L}_{\text{TB}}(\tau)=\Big{(}\mathrm{log}\frac{\mathcal{Z}\prod_{l=0}^{L-1}\mathbb{P}_{F}(s_{l+1}|s_{l})}{\mathcal{R}(s_{L})\prod_{l=0}^{L-1}\mathbb{P}_{B}(s_{l}|s_{l+1})}\Big{)}^{2}$
Here, $\mathcal{Z}$ is a global scalar that approximates $\mathcal{F}(s_{0})$ (Malkin et al., 2022).
Motivation. Two key insights motivate us to apply GFlowNets for GR: first, GR inherently involves a multi-step generation process well-suited for GFlowNet modeling. As shown in Figure 1 (b), generating the next item identifier unfolds over $L$ sequential steps: starting from an initial state $s_{0}=\{\emptyset\}$, LLMs iteratively generate tokens (e.g., $<a\_2>$, $<b\_4>$, $<c\_10>$) to form intermediate states $s_{1}=\{<a\_2>\}$, $s_{2}=\{<a\_2><b\_4>\}$, and finally reach the terminal state $s_{3}=\{<a\_2><b\_4><c\_10>\}$, which decodes to a specific item (e.g., a blue polo). This sequential structure aligns naturally with GFlowNets’ trajectory-based framework. Second, GFlowNets offer two critical advantages for GR: (a) Diverse High-Reward Generation: by linking generation probability to reward ($\mathbb{P}(\tau)\propto\mathcal{R}(s_{L})$), GFlowNets can explore and generate diverse items with high predicted rewards. Theoretical analyses (Bengio et al., 2021, 2023) and applications in domains like DNA design (Jain et al., 2022) and multi-step reasoning (Kang et al., 2025; Yu et al., 2024) validate this capability. In recommendation, this addresses exposure bias by treating unobserved interactions as potential positives rather than default negatives. Particularly, during training, GFlowNets can identify augmented items with high reward scores as potential positive and prioritize them over low-reward alternatives. (b) Off-Policy Data Utilization: GFlowNets enable off-policy learning, allowing GR frameworks to incorporate augmented interactions into training. This extends SFT, which relies solely on collected positive samples, to leverage a broader range of potential user preferences. The lower part of Figure 1 (b) illustrates these: by associating generation probabilities with rewards, GFlowNets empower GR to prioritize the unseen blue jacket (aligned with the user’s formal dressing history) over the yellow undershirt, even when both are unobserved. This contrasts with conventional methods that treat unobserved items equally, thereby enhancing diversity and mitigating exposure bias.
Implementing GFlowNets for GR involves straightforward computation of forward and backward probabilities: the forward probability $\mathbb{P}_{F}(s_{l+1}|s_{l})$ corresponds to the LLM’s token generation probability $\mathbb{P}_{\text{GR}}(t_{l+1}|U,s_{l})$; the backward probability $\mathbb{P}_{B}(s_{l}|s_{l+1})=1$ due to the sequential nature of token generation ($s_{l+1}=\{t_{i}\}_{i\leq(l+1)}=\{s_{l},t_{l+1}\}$, $s_{l+1}$ is uniquely derived from $s_{l}$). Thus, the core design challenges reduce to constructing training trajectories $\tau$ and designing the reward model $\mathcal{R}(s_{L})$. These components are critical for GFlowNets and will be addressed in subsequent sections.
3. Methodology
In this section, we detail the proposed GFlowGR framework, including its adaptive trajectory sampler, comprehensive reward model, and training/inference procedures. The advantages of GFlowGR in addressing exposure bias and negative sampling are summarized in the Supplementary Material.
The architecture of GFlowGR is depicted within the dashed block of Figure 2. Notably, GFlowGR modifies only the LLM fine-tuning stage for GR while keeping the inference pipeline unchanged. The framework integrates a collaborative model (CM) that is capable of scoring candidate items (blue dotted block) based on user interaction histories (pink dotted block) as a core component of both the trajectory sampler and reward model. The overall workflow is as follows: for a given input pair $(U,v)$, the adaptive trajectory sampler selects augmented samples $v^{\prime}$ as extended generation trajectories. GFlowNet then estimates the forward probabilities with $\mathbb{P}_{F}$ and flow values with $\mathcal{F}$ for both the original trajectory $\tau$ (corresponding to the ground-truth $v$) and augmented trajectories $\tau^{\prime}$ (corresponding to $v^{\prime}$). The reward model $\mathcal{R}$ integrates collaborative scores ($r_{c}$), augmentation signals from the trajectory sampler ($r_{a}$), and token similarities between positive and augmented item identifiers ($r_{\text{sim}}$) to compute rewards. Finally, GFlowGR computes the GFlowNet loss $\mathcal{L}_{\text{GFN}}$ for the LLM fine-tuning.
Figure 2. Framework of GFlowGR. The visualized example shows a positive sample (blue polo shirt) and an augmented sample (yellow undershirt). Forward probabilities ($\mathbb{P}_{F}$) for both trajectories are displayed (e.g., $0.9,0.8,0.83$ for $\tau$ ), while estimated flow values ($\mathcal{F}$) are omitted for brevity.
3.1. Adaptive Trajectory Sampler
The sparsity of recommendation data is a fundamental challenge, as only a minimal fraction of user-item interactions are observed (Gao et al., 2022). Consequently, constructing training trajectories solely from collected data is insufficient for GFlowGR, rendering the trajectory sampler essential to augment the sparse dataset. A naive approach to address this is through data augmentation (Dang et al., 2024) or random negative sampling (Ma et al., 2024)—for instance, randomly selecting $K$ augmented samples $\{v^{\prime}_{k}\}_{k\leq K}$ such that $v\neq v^{\prime}_{k}$ for all $k$ (Yang et al., 2020; Wang et al., 2021b). However, such randomness introduces training instability and often yields samples misaligned with the model’s evolving learning dynamics. For example, an obviously irrelevant augmented sample that was effective as a negative in early training epochs may become uninformative in later stages, as the model learns to easily distinguish it from true positives.
To address this, we propose an adaptive trajectory sampler that delivers high-quality augmented samples to GFlowGR in a curriculum-style manner. We assume that during the early training phase, LLMs struggle to distinguish between positive and augmented samples, while in later stages, overly simple samples contribute minimally to model improvement (Wang et al., 2021a). Inspired by recent work demonstrating the benefits of integrating collaborative knowledge into LLM-based recommendation systems (Lin et al., 2025a), we leverage a pretrained collaborative model (CM) to guide sample augmentation. Specifically, we score $N$ candidate augmentations $\{v^{\prime}_{n}\}_{n\leq N}$ using $\hat{y}_{CM}^{v^{\prime}_{n}}=\mathrm{CM}(v^{\prime}_{n},U)$, where $\mathrm{CM}(\cdot)$ denotes the collaborative model and $\hat{y}_{CM}^{v^{\prime}_{n}}$ represents the collaborative score of item $v^{\prime}_{n}$ to user $U$.
During initial training epochs, we prioritize sampling items with the lowest $\hat{y}_{CM}^{v_{n}}$ (easy negatives), such as the yellow undershirt in Figure 1(b) and Figure 2, which has low relevance to a user preferring blue formal attire. As training progresses, the sampler gradually incorporates harder augmented samples with higher collaborative scores—for example, the blue jacket in the visualized example, which aligns more closely with the user’s preferences but remains unobserved in historical data. This curriculum-based approach enables the sampler to adaptively select trajectories rooted in collaborative knowledge, reducing randomness and enhancing training stability by ensuring augmented samples remain challenging yet relevant at each stage of learning.
3.2. Comprehensive Reward Model
The reward model is pivotal in GFlowNets, as it directly aligns generation probabilities with task-specific objectives (Bengio et al., 2021, 2023). In recommendation contexts, prior GFlowNet reward models typically rely on observed user interactions (e.g., item ratings from 1–6 (Liu et al., 2024b, 2023)), but this approach fails for unobserved samples. Assigning zero rewards to augmented samples may ignore the latent relevance and exacerbate exposure bias (Chen et al., 2023).
To address this, we introduce a comprehensive reward model that integrates three complementary signals: augmentation status, collaborative scores, and token similarity. For simplicity, we describe the framework using a single augmented sample ($K=1$, easily extendable to multiple samples). Let $\tau=\{s_{l}\}_{l\leq L}$ denote the trajectory of the observed positive item $v$, and $\tau^{\prime}=\{s^{\prime}_{l}\}_{l\leq L}$ denote the trajectory of an augmented item $v^{\prime}$. Three aspects of reward are defined as:
  • Augmentation Signal ($r_{a}$): marks whether a sample is observed (positive) or augmented, i.e., $r_{a}=1$ for the ground-truth trajectory $\tau$ and $r^{\prime}_{a}=0$ for the augmented trajectory $\tau^{\prime}$.
  • Collaborative Score ($r_{c}$): the collaborative score from the same model, i.e., $r_{c}=\hat{y}_{CM}^{v},r^{\prime}_{c}=\hat{y}_{CM}^{v^{\prime}}$.
  • Token Similarity ($r_{\text{sim}}$): rewards partial correctness by measuring the number of shared tokens with the positive sample. Specifically, for $\tau^{\prime}$, $r^{\prime}_{\text{sim}}=\sum_{l=1}^{L}\mathbb{I}(t_{l}=t^{\prime}_{l})$, where $\mathbb{I}(\cdot)$ is an indicator function. And $r_{\text{sim}}=L$ for $\tau$. For the visualized example, $r^{\prime}_{\text{sim}}=1$, as it has one shared token with the ground truth.
We sum up these reward signals as the final reward for each trajectory, which can be formally expressed as $\mathcal{R}(\tau)=r_{a}+r_{c}+r_{\text{sim}}$.
3.3. Training and Inference
Training. The training objective of GFlowGR integrates two complementary loss components: the vanilla next-token prediction loss $\mathcal{L}_{\text{GR}}$ for positive samples, the GFlowNet training loss $\mathcal{L}_{\text{GFN}}$ over sampled trajectories. The inclusion of $\mathcal{L}_{\text{GR}}$ (Equation (1)) ensures the model retains foundational recommendation capabilities while adapting the LLM to GR tasks. For $\mathcal{L}_{\text{GFN}}$, substituting the forward probability $\mathbb{P}_{F}(s_{l+1}|s_{l})=\mathbb{P}_{\text{GR}}(t_{l+1}|U,s_{l})$ and backward probability $\mathbb{P}_{B}(s_{l}|s_{l+1})=1$ into Equation (3) and (4) yields:
(5) $\mathcal{L}_{\text{DB}}(\tau)=\sum_{l=0}^{L-1}\Big{(}\mathrm{log}\frac{\mathcal{F}(s_{l})\mathbb{P}_{\text{GR}}(t_{l+1}|U,s_{l})}{\mathcal{F}(s_{l+1})}\Big{)}^{2}$
(6) $\mathcal{L}_{\text{TB}}(\tau)=\Big{(}\mathrm{log}\frac{\mathcal{Z}\prod_{l=0}^{L-1}\mathbb{P}_{\text{GR}}(t_{l+1}|U,s_{l})}{\mathcal{R}(s_{L})}\Big{)}^{2}$
$\mathcal{F}$, and $\mathcal{Z}$ are learnable GFlowNet components, with $\mathcal{F}(s_{L})=\mathcal{R}(s_{L})$ in Equation (5). $\mathcal{L}_{\text{GFN}}$ can be instantiated using either $\mathcal{L}_{\text{DB}}$ or $\mathcal{L}_{\text{TB}}$. The complete loss function for $(U,v,\tau)$ is:
(7) $\mathcal{L}(U,v,\tau)=\mathcal{L}_{\text{GR}}(U,v)+\lambda\sum_{\kappa\in\{\tau,\tau^{\prime}\}}\mathcal{L}_{\text{GFN}}(\kappa)$
where $\kappa$ represents either the positive trajectory $\tau$ or an augmented trajectory $\tau^{\prime}$, and $\lambda$ is the hyperparameter to balance different objectives. To mitigate exposure bias, we weight each trajectory by the inverse disagreement between the augmentation signal $r_{a}$ and the collaborative score $r_{c}$:
(8) $\mathcal{L}(U,v,\tau)=\mathcal{L}_{\text{GR}}(U,v)+\lambda\sum_{\kappa\in\{\tau,\tau^{\prime}\}}\frac{1}{(r_{a}-r_{c})^{2}}\mathcal{L}_{\text{GFN}}(\kappa)$
This heuristic weighting strategy downweights examples where the collaborative model exhibits high uncertainty (large $|r_{a}-r_{c}|$), which often correspond to hard negatives or ambiguous interactions (Chen et al., 2023).
Table 1. Overall performance of baselines and GFlowGR. The boldface refers to beating all baselines. “*” indicates the statistically significant improvements (i.e., one-sided t-test with $p<0.05$) over the best baseline. For all metrics, the higher is better.
Model Finetuning Beauty Yelp
R@5 R@10 N@5 N@10 R@5 R@10 N@5 N@10
TIGER SFT 0.0354 0.0558 0.0234 0.0299 0.0204 0.0334 0.0132 0.0174
PPO 0.0020 0.0032 0.0013 0.0017 0.0004 0.0005 0.0003 0.0003
GRPO 0.0347 0.0543 0.0229 0.0291 0.0193 0.0328 0.0126 0.0169
DPO 0.0308 0.0483 0.0200 0.0256 0.0100 0.0182 0.0060 0.0087
S-DPO 0.0314 0.0492 0.0208 0.0265 0.0212 0.0333 0.0137 0.0176
SPRec 0.0312 0.0487 0.0205 0.0258 0.0103 0.0186 0.0067 0.0090
IPA 0.0306 0.0481 0.0198 0.0245 0.0101 0.0179 0.0066 0.0093
GFlowGR-DB 0.0388* 0.0629* 0.0254* 0.0336* 0.0221 0.0345 0.0145 0.0185
GFlowGR-TB 0.0407* 0.0651* 0.0266* 0.0344* 0.0256* 0.0404* 0.0166* 0.0213*
LETTER SFT 0.0344 0.0540 0.0227 0.0290 0.0205 0.0341 0.0131 0.0175
PPO 0.0128 0.0206 0.0087 0.0112 0.0002 0.0004 0.0001 0.0002
GRPO 0.0384 0.0592 0.0251 0.0317 0.0228 0.0380 0.0146 0.0195
DPO 0.0358 0.0571 0.0238 0.0307 0.0172 0.0286 0.0115 0.0151
S-DPO 0.0390 0.0607 0.0259 0.0336 0.0247 0.0389 0.0161 0.0206
SPRec 0.0370 0.0573 0.0249 0.0315 0.0190 0.0303 0.0124 0.0160
IPA 0.0367 0.0586 0.0252 0.0320 0.0197 0.0293 0.0129 0.0163
GFlowGR-DB 0.0407* 0.0630* 0.0266 0.0337 0.0252 0.0399 0.0166 0.0213
GFlowGR-TB 0.0417* 0.0632* 0.0276* 0.0345* 0.0261* 0.0413* 0.0173* 0.0222*
Inference. The inference process remains unchanged from standard GR frameworks: user prompts are constructed as before, and the fine-tuned LLM generates item identifiers sequentially. By design, GFlowGR enhances GR without altering the inference pipeline, ensuring seamless integration into existing systems.
4. Experiments
4.1. Experimental Setups
Datasets. We used two real-world datasets to evaluate GFlowGR, namely Yelp and Amazon Beauty. The data processing and split are following previous GR works (Wang et al., 2024a; Zheng et al., 2024).
Baselines and Backbones. We compare GFlowGR with three categories of LLM fine-tuning baselines: (a) vanilla SFT; (b) on-policy RLFT: PPO (Schulman et al., 2017), GRPO (Shao et al., 2024); (c) off-policy RLFT: DPO (Rafailov et al., 2023), and DPO-based algorithms modified for language-based recommendations, including S-DPO (Chen et al., [n. d.]), SPRec (Gao et al., 2025a), IPA (Deng et al., 2025). Notably, the specially designed auxiliary loss in addition to SFT is regarded as a part of the GR backbones. And we install GFlowGR to two backbones: TIGER (Rajput et al., 2023) and LETTER (Wang et al., 2024a).
Evaluation Metrics. We evaluate all methods using two metrics, namely Hit Rate (R) and Normalized Discounted Cumulative Gain (NDCG), on top-5 and top-10 recommended lists, yielding four metrics per dataset: R@5, R@10, N@5, and N@10. To ensure robust results, each test is repeated three times with different random seeds ({42, 43, 44}).
4.2. Main Results
To evaluate the effectiveness of GFlowGR, we tested its two variants (DB and TB) against various fine-tuning methods. The results are summarized in Table 1, which consistently demonstrates that GFlowGR outperforms baseline approaches. Below are detailed observations:
Comparison with SFT. Both GFlowGR variants consistently outperform SFT across all datasets and metrics, demonstrating that GFlowGR effectively enhances model capacity through augmented samples, flow-matching objectives, and integrated collaborative knowledge. Notably, GFlowGR-TB achieves more significant improvements than GFlowGR-DB, particularly on Yelp with TIGER, likely because GR’s intermediate states do not uniquely identify items, making whole-trajectory optimization (via TB’s trajectory-level balance) more effective by leveraging one-to-one trajectory-item correspondences.
Comparison with On-policy RLFT. On-policy strategies (especially reward-based PPO) fail to achieve competitive results for GR. This may stem from out-of-vocabulary item identifiers: LLMs pretrained on general language tasks struggle to explore the OOV item space without a well-defined reward model, leading to poor PPO performance. GRPO alleviates this dependency on the reward model via relative advantages. (Shao et al., 2024)
Comparison with Off-policy RLFT. Vanilla DPO (Rafailov et al., 2023) fails to stably improve upon SFT, particularly on large-scale Yelp data, due to its use of fixed and randomly sampled “Rejected” items. In contrast, SPRec (Gao et al., 2025a) (using self-play sampling) and IPA (Deng et al., 2025) (leveraging collaborative knowledge) enhance DPO through dynamic, guided negative sampling. Among DPO-based methods, S-DPO (Chen et al., [n. d.]), which is capable of incorporating multiple negative samples, achieves the best results and competes closely with GFlowGR-DB.
Figure 3. Parameter study on $K$ and $\lambda$ with TIGER on Beauty. The red dashed line denotes the performance of SFT, which falls outside the plotted range for the figures corresponding to $K$.
4.3. Parameter Analysis
In this section, we analyze GFlowGR ’s parameter sensitivity, focusing on the number of augmented trajectories $K\in\{0,1,3,5\}$ and the balancing weight $\lambda\in\{0.01,0.1,1,10,100\}$ in Equation (8). Figure 3 visualizes results for TIGER on the Beauty dataset. As $K$ and $\lambda$ increase, performance follows a reverse ‘V’ curve: For $K$, small values (e.g., $K=0$) fail to provide sufficient diversity for sparse data, while large $K=5$ introduces noisy augmented samples that mislead training. For $\lambda$, a moderate weight of $1.0$ optimizes $\mathcal{L}_{\text{GFN}}$, balancing flow-matching objectives with the core recommendation task. Overemphasizing $\mathcal{L}_{\text{GFN}}$ (e.g., $\lambda\geq 10$) degrades GR performance, highlighting the critical role of SFT in maintaining basic recommendation capabilities.
4.4. Ablation Status
In this section, we validate the critical components of GFlowGR: the curriculum-style adaptive trajectory sampler, comprehensive reward model, and uncertainty-based trajectory reweighting. Table 2 presents ablation results for: all reward signal combinations (six variants, first block of the table, where the name of the variant denotes used reward dignals), removal of adaptive sampling (w/o Ada), omission of Augmented Trajectories (w/o Traj), and removal of reweighting (w/o Wt). Key findings include: (a) Loss of any reward signal (augmentation signals, collaborative scores, or token similarities) degrades performance, confirming the necessity of the comprehensive reward model. Notably, $r_{a}$ has the most significant impact, while $r_{c}$ contributes marginally due to CM producing less accurate predictions than GR (Rajput et al., 2023; Wang et al., 2024a), leaving to noisy $r_{c}$ signals. (b) Removing adaptive sampling or using random sampling reduces effectiveness due to unstable training or low-quality augmented samples. (c) Uncertainty-based reweighting contributes to alleviating exposure bias and improving model robustness, validating its role in stabilizing training.
Table 2. Ablation study with TIGER on Beauty.
Metrics $r_{a}$ $r_{c}$ $r_{\text{sim}}$ $r_{a}+r_{c}$ $r_{a}+r_{\text{sim}}$ $r_{c}+r_{\text{sim}}$ w/o Ada w/o Traj w/o Wt GFlowGR
R@5 0.0403 0.0399 0.0404 0.0399 0.0406 0.0381 0.0386 0.0397 0.0388 0.0407
R@10 0.0641 0.0631 0.0634 0.0629 0.0648 0.0600 0.0600 0.0617 0.0615 0.0651
N@5 0.0261 0.0263 0.0260 0.0264 0.0260 0.0253 0.0257 0.0261 0.0254 0.0266
N@10 0.0341 0.0338 0.0343 0.0337 0.0344 0.0323 0.0326 0.0332 0.0327 0.0344
4.5. In-depth Analysis
Case Study. We visualize a case in Figure 4 to study GFlowGR’s potential in addressing exposure bias. The example shows GFlowGR recommending a diverse set of items, including ‘Shampoo’, ‘Scrub’, and ‘Cleanser’ with the ‘Tomatoes’ flavor (bold green), which is a new taste for the user. While SFT focuses on ‘Carrots’-flavored ‘Shampoo’ and ‘Conditioner’ from the same brand, which the user recently purchased (bold orange and red). This suggests GFlowGR alleviates exposure bias by exploring diverse trajectories, generating exploratory results with different flavors and product.
Diversity. For this experiment, we analyze diversity in top-10 recommendations from GFlowGR and SFT across two dimensions. (a) Generation Distribution: We compare models’ item recommendation probabilities against a normal distribution using KL divergence (rather than a uniform distribution because item popularities are inherently different for recommendations (Chen et al., 2023)). GFlowGR achieves a closer match (KL=3.13 vs. SFT’s 3.24). In addition, the generation probability over all items has a 2.2% smaller standard deviation with GFlowGR, indicating a smoother, more diverse distribution. (b) Collaborative Scores: We calculate the collaborative score for the recommended items. The result shows that GFlowGR’s recommended items have 0.4% lower average collaborative scores but a broader range (0.45 vs. SFT’s 0.42) and 3.6% higher standard deviation, showing it prioritizes non-mainstream items while offering greater score diversity. These results demonstrate GFlowGR’s enhanced ability to balance relevance and novelty compared to SFT.
Figure 4. Case study on user-2089 from Beauty with TIGER: We visualized three items from the top-10 list of each method for the brand ‘Yes to’, where the top-1 result of GFlowGR is correct.
5. Related Work
Generative Recommendations. Existing GR research can be categorized into three directions: pipeline design, item tokenizer innovations, and LLM decoding enhancement. For pipeline design, TIGER (Rajput et al., 2023) pioneered out-of-vocabulary item tokenization and GR workflows, while LC-Rec (Zheng et al., 2024) and ColaRec (Wang et al., 2024b) enhance alignment via tasks or graph collaborative models, with extensions to multi-behavior scenarios (Liu et al., 2024a). The second category focuses on plug-in tokenizers (e.g., LETTER (Wang et al., 2024a) for collaborative embeddings) to boost arbitrary GR frameworks. The third category optimizes decoding (Lin et al., 2025c; Ding et al., 2024). However, LLM fine-tuning for GR remains underexplored, with DPO-based IPA (Deng et al., 2025) as a rare example. The proposed GFlowGR is the first to study the exposure bias and diversity.
Generative Flow Network Applications. GFlowNets are designed for compositional generation to produce diverse, high-reward outputs (Bengio et al., 2021, 2023), with applications spanning biology (Jain et al., 2022), scientific discovery (Jain et al., 2023), and multi-step reasoning (Yu et al., 2024; Kang et al., 2025). While prior work has applied GFlowNets to discriminative recommendation tasks like sequential recommendation (Liu et al., 2024b) and list-wise recommendation (Liu et al., 2023), as well as language-based recommendations (Gao et al., 2025b), we are the first to explore GFlowNets for generative recommendations.
6. Conclusion
In this paper, we propose GFlowGR, a GFlowNets-based fine-tuning framework for GR. Specifically, we design an adaptive trajectory sampler that leverages collaborative knowledge to select augmented trajectories, enabling the model to explore diverse item spaces. We also introduce a comprehensive reward model that integrates multi-faceted signals, including augmentation status, collaborative scores, and token similarities, to provide nuanced feedback for trajectory evaluation. During fine-tuning, GFlowGR extends the vanilla SFT loss by incorporating flow-matching objectives, with trajectory-specific losses reweighted based on uncertainty. By adaptive sampling, heuristic reweighting, and flow-matching targets, GFlowGR enhances recommendation diversity and addresses the exposure bias challenge inherent in GR frameworks. Extensive experimental results validate the effectiveness of our approach across multiple datasets, backbones, and metrics.

### v2

@@
 Abstract.
-Generative recommendations (GR), which usually include item tokenizers and generative Large Language Models (LLMs), have demonstrated remarkable success across a wide range of scenarios.
-The majority of existing research efforts primarily concentrate on developing powerful item tokenizers or advancing LLM decoding strategies to attain superior performance.
-However, the critical fine-tuning step in GR frameworks, which is essential for adapting LLMs to recommendation data, remains largely unexplored.
-Current approaches predominantly rely on either the next-token prediction loss of supervised fine-tuning (SFT) or recommendation-specific direct preference optimization (DPO) strategies.
-Both methods ignore the exploration of possible positive unobserved samples, which is commonly referred to as the exposure bias problem.
-To mitigate this problem, this paper treats the GR as a multi-step generation task and constructs a GFlowNets-based fine-tuning framework (GFlowGR).
-The proposed framework integrates collaborative knowledge from traditional recommender systems to create an adaptive trajectory sampler and a comprehensive reward model.
-Leveraging the diverse generation property of GFlowNets, along with sampling and heuristic weighting techniques, GFlowGR emerges as a promising approach to mitigate the exposure bias problem.
-Extensive empirical results on two real-world datasets and with two different GR backbones highlight the effectiveness and robustness of GFlowGR.
+Generative recommendation (GR) systems, which leverage Large Language Models (LLMs) to produce recommendations, have demonstrated remarkable success in industrial applications.
+However, despite significant research into item tokenization and decoding strategies, a critical gap persists between the standard training paradigm and the demands of real-world deployment.
+Existing methods often fail to effectively leverage the rich, value-aware signals in user interaction data.
+They typically fine-tune models on a single ground-truth item in a value-unaware manner, while even reward-based approaches often neglect crucial token-level supervision, ultimately limiting performance.
+To address these limitations, we formulate the GR task as a multi-step generation problem and introduce GFlowGR, a novel fine-tuning framework based on Generative Flow Networks (GFlowNet).
+The modular design of GFlowGR features three core components: a trajectory sampler, a reward model, and the GFlowNet loss.
+We present the various configurations of GFlowGR for diverse practical scenarios.
+Extensive experiments on three real-world datasets using two GR backbones demonstrate the effectiveness of GFlowGR.
+For real-world deployment, GFlowGR is live across all of Taobao’s advertising businesses, driving a 1% relative increase in billion-level annual revenue since 2025 May.
 Recommender Systems; Generative Recommendations
@@
 * Xiangyu Zhao is corresponding author.
-copyright: acmcopyrightjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06
+copyright: acmcopyrightjournalyear: 2026doi: XXXXXXX.XXXXXXXconference: Proceedings of the ACM Web Conference 2026; April 13–17, 2026; Dubai, UAEisbn: 978-1-4503-XXXX-X/18/06
 1. Introduction
@@
 As a result, they are increasingly utilized across various domains, including healthcare, education, and science (Naveed et al., 2023), as well as for information retrieval and recommendations in numerous online services, such as E-commerce (Li et al., 2024a; Zhai, 2024).
-Notably, the impressive generative capabilities of LLMs have spurred the development of generative recommendation (GR) frameworks (Rajput et al., 2023; Huang et al., 2024; Hua et al., 2023).
+Notably, the impressive generative capabilities of LLMs have spurred the research on generative recommendation (GR) frameworks (Rajput et al., 2023; Huang et al., 2024; Hua et al., 2023), as well as the increasing deployment in industry (Zhou et al., 2025; Zheng et al., 2025b; Guo et al., 2025; Wei et al., 2025).
 As visualized in Figure 1 (a), these frameworks typically consist of three key components (Li et al., 2024b): item tokenizers, user prompts, and generative LLMs.
-Meanwhile, the training paradigms usually consist of two stages: optimizing the item tokenizer to represent items with several tokens and fine-tuning the LLMs for recommendation objectives like next-item prediction.
-Existing research on GR has primarily concentrated on enhancing the item tokenizer (Zheng et al., 2025; Liu et al., 2024c; Lin et al., 2025b; Wang et al., 2024a; Zhu et al., 2024) and the associated optimization processes, designing user prompts (Liu et al., 2024a; Wang et al., 2024b), and advancing generative LLMs (Lin et al., 2025c; Ding et al., 2024).
-However, there has been little focus on innovating the fine-tuning step of LLMs.
-While this step is crucial for GR frameworks: the semantics of recommendations can significantly differ from those of LLMs pretrained for general language purposes (Lin et al., 2025a; Wu et al., 2024).
-For instance, terms like “Comment” represent specific user behaviors within recommendation data, even though they are commonly used verbs in general language tasks.
-Moreover, the item identifier in GR may consist of out-of-vocabulary tokens for pretrained LLMs (Huang et al., 2024; Wang et al., 2024a; Rajput et al., 2023), which are unlikely to be generated by zero-shot LLMs.
-Current LLM fine-tuning techniques applied for GR can be generally divided into two categories (Wu et al., 2025; Wang et al., [n.
-d.]): supervised fine-tuning (SFT) and reinforcement learning-based fine-tuning (RLFT).
-The former optimizes LLMs with the next token prediction loss (Brown et al., 2020; Radford et al., 2019; Hu et al., 2022) and handcrafted auxiliary alignment loss (Liu et al., 2024a; Zheng et al., 2024; Wang et al., 2024b), which focuses solely on positive samples within the dataset.
-On the one hand, SFT does not incorporate any negative samples for comparison with the positive ones, leaving a gap with the ranking objective of recommender systems (Chen et al., [n.
-d.]).
-On the other hand, SFT cannot explore potential positive samples to extend the collected data, which is always sparse (Gao et al., 2022).
-With the exploration ability, RLFT methods like proximal policy optimization (PPO) (Schulman et al., 2017), and direct preference optimization (DPO) (Rafailov et al., 2023) can improve the model capacity on the basis of SFT by achieving the maximum reward.
-For example, recent DPO-based methods facilitate LLM fine-tuning by integrating augmented negative samples for GR (Chen et al., [n.
-d.]; Bai et al., 2024; Liao et al., 2024; Gao et al., 2025a; Deng et al., 2025), enabling LLMs to rank candidate items.
-However, both solutions cannot match the diverse nature of recommender systems that users may prefer multiple items, while only one among them can be observed.
-SFT regards the unobserved interactions as negative, and RLFT methods predefine the largest reward to the observed samples.
-This reliance on the collected data that fails to fully reveal the real user preference can lead to biased user preferences, known as exposure bias (Chen et al., 2023).
-Therefore, it calls for an LLM fine-tuning solution that can explore potential positive samples and leverage augmented samples to achieve diverse results.
-Recent advancements in Generative Flow Networks (GFlowNets) present a promising solution: by training LLMs to generate token sequences with probabilities proportional to a specific reward distribution (Bengio et al., 2023), GFlowNets can identify potential positive candidates and thus extend the recommendation diversity.
+Despite significant advances in item tokenization (Zheng et al., 2025a; Liu et al., 2024c; Lin et al., 2025b; Wang et al., 2024a; Zhu et al., 2024), user prompting, and inference optimization (Lin et al., 2025c; Ding et al., 2024), a critical bottleneck for deploying industrial-scale GR systems remains the LLM finetuning process.
+The standard paradigm, Supervised Fine-Tuning (SFT) (Rajput et al., 2023), is fundamentally misaligned with the core objective of production systems.
+These systems, whether deployed for retrieval (Yang et al., 2025) or as end-to-end “OneModels” (Zhou et al., 2025), are tasked with generating a diverse and high-quality set of recommendations, where items may have different values (e.g., interaction type for training and recommendation scores for prediction).
+However, SFT’s sequence-to-sequence objective is problematic for two key reasons.
+First, it trains the model to predict only a single ground-truth item per instance, ignoring the rest of the relevant item set.
+Second, it often fails to model the varying value across interacted items.
+These create a significant gap between offline training and real-world deployment objectives for online serving.
+To bridge this gap, recent work has turned to Reinforcement Learning (RL) to better exploit the full, value-aware item set (Chen et al., 2025; Zhou et al., 2025).
+For example, Chen et al.
+(2025) adapt Direct Preference Optimization (DPO) (Rafailov et al., 2023) by constructing preference pairs from the user’s interaction history; items with higher scores (e.g., from longer dwell times or clicks) are treated as the “chosen” examples, while those with lower scores become the “rejected” ones.
+Similarly, OneRec-V2 (Deng et al., 2025) modifies the group-wise learning objective GRPO (Shao et al., 2024) to leverage the entire set of real-world user interactions.
+A common limitation of these methods, however, is their reliance on outcome-based rewards.
+They assign rewards at the item level but fail to supervise the underlying token-level generation process.
+Since items in GR systems are composed of multiple tokens (Figure 1 (b)), this lack of process-level supervision represents a critical performance bottleneck (Gao et al., 2025b).
+Therefore, it calls for an effective fine-tuning solution that not only learn from a value-aware item set but also provide direct supervision over the token-by-token generation process.
+Recent advancements in Generative Flow Networks (GFlowNets) present a promising solution: Organizing the item set as generation trajectories, assigning item value as the termination reward, GFlowNets can attribute flow values as the token-level reward, which also provide the additional benefit of generating token sequences with probabilities proportional to the reward distribution (Bengio et al., 2023), enhancing the generation diversity, which is also a practical demand of industrial recommendation scenarios.
 This capability is supported by recent studies demonstrating the application of GFlowNets for multi-step reasoning tasks (Kang et al., 2025; Takase et al., 2024; Yu et al., 2024) and compositional generation (Jain et al., 2022).
-Generative recommendation tasks are inherently similar to these tasks; as illustrated in Figure 1 (b), recommending an item involves multiple generation steps of a single token.However, applying GFlowNets for GR still faces significant challenges: how to organize training trajectories (samples) and establish a reward model to assess the terminal state for GFlowNets.
-Specifically, the collected data consists solely of positive interaction logs from users, resulting in sparse training trajectories for GFlowNets.
-This scenario necessitates high-quality augmented trajectories for robust training.
-Furthermore, merely treating these augmented trajectories as negative samples and assigning them low rewards is inadequate for overcoming the exposure bias problem, as these augmented trajectories are not consistently negative.
-To address these limitations, this paper constructs the GFlowNets-based fine-tuning paradigm for GR (GFlowGR).
-Drawing inspiration from prior insights on aligning LLMs with collaborative recommendation models (Hong et al., [n.
-d.]; Yang et al., 2024; Lin et al., 2025a), we incorporate collaborative scores when designing the trajectory sampling and reward model for GFlowGR.
-Specifically, we develop an adaptive trajectory sampler that generates easy-to-hard augmented trajectories, thereby implementing a curriculum-style training approach (Wang et al., 2021a) to consistently provide high-quality training data.
-Additionally, we construct a multi-faceted reward system that thoroughly considers augmentation signals, collaborative scores, and semantic similarities to effectively differentiate possibly preferred items from the sampled unobserved trajectories.
-Finally, to further mitigate the exposure bias problem, we assign confidence weights to each sample based on the agreement between augmentation signals and collaborative scores.
+Generative recommendation tasks are inherently similar to these tasks; as illustrated in Figure 1 (b), recommending an item involves multiple generation steps of a single token.
+Despite the recent successes in deploying GFlowNets for traditional recommendations (Liu et al., 2023, 2024b) or text-based LLM recommendations (Gao et al., 2025b), there still lack a conceptual GFlowNets framework for GR systems.
+Consequently, this paper constructs the GFlowNet-based fine-tuning paradigm for GR (GFlowGR).
+We suggest three key components for GFlowGR: a trajectory sampler to structure training data, a reward model to estimate item value, and a GFlowNet training objective.
+We detail the conceptual implementation of these components and show how they assemble into a cohesive system.
+Designed for practical adoption, GFlowGR can cooperate with existing recommendation models, and serve as a powerful supplement to standard SFT.
+Since May 2025, GFlowGR has been deployed in Taobao’s production GR system, serving hundreds of millions of daily active users and driving billions of dollars in annual advertising revenue.
 We summarize the major contributions of this paper as follows:
-  • We formulate GR as a multi-step generation process and develop a GFlowNets-based fine-tuning framework, GFlowGR.
-This framework helps mitigate the exposure bias problem by learning generation probabilities that are proportional to the reward.
-  • We implement an adaptive trajectory sampler to select high-quality training samples for GFlowGR and establish a comprehensive reward model to consider various reward signals effectively.
-  • Our extensive experimental results, obtained from two public benchmarks using two backbone GR models, demonstrate the effectiveness and versatility of GFlowGR.
-Figure 1. Overview of GR and GFlowNets, with the length of item identifiers $L=3$.
+  • We formulate GR as a multi-step generation problem and introduce GFlowGR, a new fine-tuning paradigm based on GFlowNets that provides token-level supervision.
+  • We present the modular architecture of GFlowGR, consisting of a trajectory sampler, reward model, and GFlowNet loss.
+We demonstrate its flexibility by instantiating these components with various configurations.
+  • We show through extensive experiments on three public datasets and two backbone models that GFlowGR consistently and significantly outperforms existing methods.
+  • We detail the successful deployment of GFlowGR in Taobao’s production system, where it has been serving hundreds of millions of daily active users and driving billions of annual revenue.
+Figure 1.
+Overview of GR and GFlowNets, with the length of item identifiers $L=3$.
+(a) A process for constucting the GR framwork.
+(b) Generative flow structure is exemplified with a set of user interaction, which typically includes positive feedback (“Clicked”), impression (“Presented”), and unexposed items (“unpresented”).
 2. Preliminary
@@
 The framework first constructs an item tokenizer using item features to decompose items $v$ into token sequences $\{t_{l}\}_{l<L}$, serving as item identifiers.
-The lower portion of the figure exemplifies that a skyblue shirt is encoded as the token sequence $<a\_1><b\_4><c\_2>$ based on its title and descriptions.
-To infer the user preference, GR frameworks assemble user prompts (LLM inputs) by integrating task descriptions, historical user interactions, and profile information (e.g., age, gender denoted as $<\underline{X}>$).
+The lower portion of the figure exemplifies that a skyblue shirt is encoded as the token sequence <a_1><b_4><c_2><a\_1><b\_4><c\_2> based on its title and descriptions.
+To infer the user preference, GR frameworks assemble user prompts (LLM inputs) by integrating task descriptions, historical user interactions, and profile information (e.g., age, gender denoted as $\underline{X}>$).
 These prompts, exemplified by tokenized interaction histories (blue tokens for jackets, shirts, pants) and profiles, are fed into LLMs to generate token sequences representing the most probable next item.
-As shown in the upper portion of Figure 1 (a), this process produces $<a\_2><b\_4><c\_10>$, decoding to a blue polo shirt as the recommended item.
+As shown in the upper portion of Figure 1 (a), this process produces <a_2><b_4><c_10><a\_2><b\_4><c\_10>, decoding to a blue polo shirt as the recommended item.
 To achieve the recommendation goal, the training paradigm of GR comprises two stages: optimizing the item tokenizer ($\mathcal{L}_{T}$) and the generative LLMs ($\mathcal{L}_{\text{GR}}$), respectively.
@@
 Studies enhancing the tokenizer often incorporate auxiliary losses into $\mathcal{L}_{T}$, such as diversity (Wang et al., 2024a) and collaborative-constrained losses (Lin et al., 2025b).
-$\mathcal{L}_{\text{GR}}$ typically adopts the next-token prediction loss (Deng et al., 2025; Rajput et al., 2023; Wang et al., 2024a; Liu et al., 2024a)::::
+$\mathcal{L}_{\text{GR}}$ typically adopts the next-token prediction loss (Deng et al., 2025; Rajput et al., 2023; Wang et al., 2024a; Liu et al., 2024a)::
 (1) $\mathcal{L}_{\text{GR}}(U,v)=-\sum_{l=1}^{L}\mathrm{log}\mathbb{P}_{\text{GR}}(t_{l}|U,\{t_{i}\}_{i<l})$
@@
 Two key insights motivate us to apply GFlowNets for GR: first, GR inherently involves a multi-step generation process well-suited for GFlowNet modeling.
-As shown in Figure 1 (b), generating the next item identifier unfolds over $L$ sequential steps: starting from an initial state $s_{0}=\{\emptyset\}$, LLMs iteratively generate tokens (e.g., $<a\_2>$, $<b\_4>$, $<c\_10>$) to form intermediate states $s_{1}=\{<a\_2>\}$, $s_{2}=\{<a\_2><b\_4>\}$, and finally reach the terminal state $s_{3}=\{<a\_2><b\_4><c\_10>\}$, which decodes to a specific item (e.g., a blue polo).
+As shown in Figure 1 (b), generating the next item identifier unfolds over $L$ sequential steps: starting from an initial state $s_{0}=\{\emptyset\}$, LLMs iteratively generate tokens (e.g., <a_2><a\_2>, <b_4><b\_4>, <c_10><c\_10>) to form intermediate states $s_{1}=\{<a\_2>\}$, $s_{2}=\{<a\_2><b\_4>\}$, and finally reach the terminal state $s_{3}=\{<a\_2><b\_4><c\_10>\}$, which decodes to a specific item (e.g., a blue polo).
 This sequential structure aligns naturally with GFlowNets’ trajectory-based framework.
-Second, GFlowNets offer two critical advantages for GR: (a) Diverse High-Reward Generation: by linking generation probability to reward ($\mathbb{P}(\tau)\propto\mathcal{R}(s_{L})$), GFlowNets can explore and generate diverse items with high predicted rewards.
+Second, GFlowNets offer two critical advantages for GR: (a) Item Value Perception: By linking generation probability to reward ($\mathbb{P}(\tau)\propto\mathcal{R}(s_{L})$), GFlowNets can explore and generate diverse items with high rewards (i.e., values).
 Theoretical analyses (Bengio et al., 2021, 2023) and applications in domains like DNA design (Jain et al., 2022) and multi-step reasoning (Kang et al., 2025; Yu et al., 2024) validate this capability.
-In recommendation, this addresses exposure bias by treating unobserved interactions as potential positives rather than default negatives.
-Particularly, during training, GFlowNets can identify augmented items with high reward scores as potential positive and prioritize them over low-reward alternatives.
-(b) Off-Policy Data Utilization: GFlowNets enable off-policy learning, allowing GR frameworks to incorporate augmented interactions into training.
-This extends SFT, which relies solely on collected positive samples, to leverage a broader range of potential user preferences.
-The lower part of Figure 1 (b) illustrates these: by associating generation probabilities with rewards, GFlowNets empower GR to prioritize the unseen blue jacket (aligned with the user’s formal dressing history) over the yellow undershirt, even when both are unobserved.
-This contrasts with conventional methods that treat unobserved items equally, thereby enhancing diversity and mitigating exposure bias.
+This addresses the need for ranking ability in recommendations: preferred items are required to possess a higher rank than low-value items.
+(b) Full Set Utilization: GFlowNets enable the learning from a sort of trajectories, allowing GR frameworks to incorporate all interactions into the single training step.
+This extends SFT, which relies solely on the single positive sample, to leverage a broader range of user interactions.
+The lower part of Figure 1 (b) illustrates these: by associating generation probabilities with rewards/values and including all interactions (clicked, presented but not clicked, and unpresented), GFlowNets empower GR to prioritize the presented blue jacket (have a larger probability to be clicked) over the yellow undershirt with a lower reward for unpresented items.
 Implementing GFlowNets for GR involves straightforward computation of forward and backward probabilities: the forward probability $\mathbb{P}_{F}(s_{l+1}|s_{l})$ corresponds to the LLM’s token generation probability $\mathbb{P}_{\text{GR}}(t_{l+1}|U,s_{l})$; the backward probability $\mathbb{P}_{B}(s_{l}|s_{l+1})=1$ due to the sequential nature of token generation ($s_{l+1}=\{t_{i}\}_{i\leq(l+1)}=\{s_{l},t_{l+1}\}$, $s_{l+1}$ is uniquely derived from $s_{l}$).
@@
 3. Methodology
-In this section, we detail the proposed GFlowGR framework, including its adaptive trajectory sampler, comprehensive reward model, and training/inference procedures.
-The advantages of GFlowGR in addressing exposure bias and negative sampling are summarized in the Supplementary Material.
-The architecture of GFlowGR is depicted within the dashed block of Figure 2.
+In this section, we detail the proposed GFlowGR framework, including the organization of trajectory sampler, reward model, and GFN losses.
+We also present the possible integration of existing collaborative models and elaborate the training and inference procedures.
+The architecture of GFlowGR is depicted within Figure 2.
 Notably, GFlowGR modifies only the LLM fine-tuning stage for GR while keeping the inference pipeline unchanged.
-The framework integrates a collaborative model (CM) that is capable of scoring candidate items (blue dotted block) based on user interaction histories (pink dotted block) as a core component of both the trajectory sampler and reward model.
-The overall workflow is as follows: for a given input pair $(U,v)$, the adaptive trajectory sampler selects augmented samples $v^{\prime}$ as extended generation trajectories.
-GFlowNet then estimates the forward probabilities with $\mathbb{P}_{F}$ and flow values with $\mathcal{F}$ for both the original trajectory $\tau$ (corresponding to the ground-truth $v$) and augmented trajectories $\tau^{\prime}$ (corresponding to $v^{\prime}$).
-The reward model $\mathcal{R}$ integrates collaborative scores ($r_{c}$), augmentation signals from the trajectory sampler ($r_{a}$), and token similarities between positive and augmented item identifiers ($r_{\text{sim}}$) to compute rewards.
+The overall workflow is as follows: for a given input pair $(U,v)$, where $v$ denotes the single item with the positive feedback [^1] In fact, the number of positive items is usually larger than one in the session.
+We discuss with the single positive situation for clarity..
+Trajectory sampler selects augmented samples or loads the collected system logs to extend the item set as $\{v_{n}\}_{n\leq N})$, where $v_{1}=v$.
+GFlowNet then estimates the forward probabilities with $\mathbb{P}_{F}$ and flow values with $\mathcal{F}$ for all trajectories $\tau_{n}$ (corresponding to $v_{n}$).
+The reward model then assigns a reward $\mathcal{R}$ for each trajectory.
 Finally, GFlowGR computes the GFlowNet loss $\mathcal{L}_{\text{GFN}}$ for the LLM fine-tuning.
+It is also possible for the framework to integrate a existing collaborative model (CM): for the candidate item $v_{n}$, CM is capable of providing the score based on user interaction histories for both the trajectory sampler and reward model.
 Figure 2.
 Framework of GFlowGR.
-The visualized example shows a positive sample (blue polo shirt) and an augmented sample (yellow undershirt).
-Forward probabilities ($\mathbb{P}_{F}$) for both trajectories are displayed (e.g., $0.9,0.8,0.83$ for $\tau$ ), while estimated flow values ($\mathcal{F}$) are omitted for brevity.
-3.1. Adaptive Trajectory Sampler
+The visualized example shows a the learning on a NN-item set illustrated with two items among the set, including a positive sample (blue polo shirt) and an unpresented sample (yellow undershirt).
+Forward probabilities ($\mathbb{P}_{F}$) for both trajectories are displayed (e.g., $0.9,0.8,0.83$ for $\tau_{1}$ ), while estimated flow values ($\mathcal{F}$) are omitted for brevity.
+The bluesky shaded area presents a example of incorporating collaborative models (CM).
+3.1. Trajectory Sampler
+This section details the construction of the trajectory sampler, which produces the learning trajectories for a given user-interaction pair $(U,v)$.
+There are four methods for constructing the trajectory set: interaction log, random sampling, CM-based sampling, and on-policy sampling.
+1.
+Interaction log.
+For the industrial system, it is straightforward to cutoff the user interaction log for the trajectory set, which typically contains more than $N$ samples.
+Specifically, with this solution, the trajectory sampler simply finds the session for $(U,v)$ and selects the N1N-1 samples to formulate the item set.
+2.
+Random sampling.
 The sparsity of recommendation data is a fundamental challenge, as only a minimal fraction of user-item interactions are observed (Gao et al., 2022).
-Consequently, constructing training trajectories solely from collected data is insufficient for GFlowGR, rendering the trajectory sampler essential to augment the sparse dataset.
-A naive approach to address this is through data augmentation (Dang et al., 2024) or random negative sampling (Ma et al., 2024)—for instance, randomly selecting $K$ augmented samples $\{v^{\prime}_{k}\}_{k\leq K}$ such that $v\neq v^{\prime}_{k}$ for all $k$ (Yang et al., 2020; Wang et al., 2021b).
+Consequently, constructing training trajectories solely from collected data is sometimes insufficient for GFlowGR, rendering the trajectory sampler essential to augment the sparse dataset.
+Consequently, we can also conduct random negative sampling (Ma et al., 2024) for the trajectory sampler.
+For instance, randomly selecting N1N-1 augmented samples $\{v_{n}\}_{2\leq n\leq N}$ from the item set such that $v_{m}\neq v_{n}$ for all $1\leq m,n\leq N,m\neq n$ (Yang et al., 2020; Wang et al., 2021b).
 However, such randomness introduces training instability and often yields samples misaligned with the model’s evolving learning dynamics.
 For example, an obviously irrelevant augmented sample that was effective as a negative in early training epochs may become uninformative in later stages, as the model learns to easily distinguish it from true positives.
-To address this, we propose an adaptive trajectory sampler that delivers high-quality augmented samples to GFlowGR in a curriculum-style manner.
+To address this, we propose two adaptive trajectory samplers that delivers high-quality augmented samples to GFlowGR in a curriculum-style manner.
 We assume that during the early training phase, LLMs struggle to distinguish between positive and augmented samples, while in later stages, overly simple samples contribute minimally to model improvement (Wang et al., 2021a).
+3.
+CM-based sampling.
 Inspired by recent work demonstrating the benefits of integrating collaborative knowledge into LLM-based recommendation systems (Lin et al., 2025a), we leverage a pretrained collaborative model (CM) to guide sample augmentation.
-Specifically, we score $N$ candidate augmentations $\{v^{\prime}_{n}\}_{n\leq N}$ using $\hat{y}_{CM}^{v^{\prime}_{n}}=\mathrm{CM}(v^{\prime}_{n},U)$, where $\mathrm{CM}(\cdot)$ denotes the collaborative model and $\hat{y}_{CM}^{v^{\prime}_{n}}$ represents the collaborative score of item $v^{\prime}_{n}$ to user $U$.
-During initial training epochs, we prioritize sampling items with the lowest $\hat{y}_{CM}^{v_{n}}$ (easy negatives), such as the yellow undershirt in Figure 1(b) and Figure 2, which has low relevance to a user preferring blue formal attire.
+Specifically, we score 2N12N-1 candidate augmentations $\{v_{n}\}_{2\leq n\leq 2N}$ using $\hat{y}_{CM}^{v_{n}}=\mathrm{CM}(v_{n},U)$, where $\mathrm{CM}(\cdot)$ denotes the collaborative model and $\hat{y}_{CM}^{v_{n}}$ represents the collaborative score of item $v_{n}$ to user $U$.
+During initial training epochs, we prioritize sampling N1N-1 items with the lowest $\hat{y}_{CM}^{v_{n}}$ (easy negatives), such as the yellow undershirt in Figure 1(b) and Figure 2, which has low relevance to a user preferring blue formal attire.
 As training progresses, the sampler gradually incorporates harder augmented samples with higher collaborative scores—for example, the blue jacket in the visualized example, which aligns more closely with the user’s preferences but remains unobserved in historical data.
-This curriculum-based approach enables the sampler to adaptively select trajectories rooted in collaborative knowledge, reducing randomness and enhancing training stability by ensuring augmented samples remain challenging yet relevant at each stage of learning.
-3.2. Comprehensive Reward Model
+4.
+On-policy sampling.
+An instant alternative to CM-based sampling is the on-policy sampling (Deng et al., 2025; Gao et al., 2025a).
+Specifically, the trajectory sampler can feed the user prompt to current LLM to generate top-(N1)(N-1) negatives for each $U$ with beam search.
+These curriculum-based approach enables the sampler to adaptively select trajectories, reducing randomness and enhancing training stability by ensuring augmented samples remain challenging yet relevant at each stage of learning.
+Algorithm 1 Train and inference process of GFlowGR with CM.
+1:Indicate the tokenizer.
+2:Indicate the LLMs.
+3:Indicate the collaborative model ($\mathrm{CM}$).
+Train Process
+5:Train the tokenzier with $\mathcal{L}_{T}$ and represent items as identifiers with the trained tokenizer.
+6:Indicate the hyperparameters for the training, including the number of augmented trajectories (NN), the weight of GFlowNets loss ($\lambda$).
+7:for a batch of samples $(U,v)$ do
+8:  Construct the user prompt with item tokenizer
+9:  Get the trajectory $\tau_{1}$ for the ground-truth $v$.
+10:  Generate N1N-1 augmented trajectories $\{\tau_{n}\}_{2\leq n\leq N}$ with the trajectory sampler with one of the strategy described in Section 3.1.
+11:  Calculate the step-wise forward probability $\mathbb{P}_{\text{GR}}(t_{l+1}|U,s_{l})$ of LLM for all trajectories.
+12:  Calculate the step-wise flow $\mathcal{F}(s_{l})$ with the flow estimator.
+13:  Calculate the reward $\mathcal{R}(s_{L})$ for the terminal state.
+14:  Calculate $\mathcal{L}_{\text{GFN}}$ with Equation (5) or Equation (6).
+15:  Calculate the SFT loss $\mathcal{L}_{\text{GR}}$ based on Equation (1).
+16:  Sum the SFT loss and GFlowNets loss for the final loss $\mathcal{L}$ as in Equation (7).
+17:  Update LLM parameters.
+18:end for
+Inference Process
+14:Load the LLM checkpoint and item tokenizer.
+15:for arbitary user $\mathcal{U}$ do
+16:  Construct the user prompt with item tokenizer.
+17:  Generate $L$ tokens as the identifier of the next item.
+18:  Decode the identifier and return the predicted item.
+19:end for
+3.2. Reward Model
 The reward model is pivotal in GFlowNets, as it directly aligns generation probabilities with task-specific objectives (Bengio et al., 2021, 2023).
@@
 Assigning zero rewards to augmented samples may ignore the latent relevance and exacerbate exposure bias (Chen et al., 2023).
-To address this, we introduce a comprehensive reward model that integrates three complementary signals: augmentation status, collaborative scores, and token similarity.
-For simplicity, we describe the framework using a single augmented sample ($K=1$, easily extendable to multiple samples).
+To address this, we introduce a comprehensive reward model that can integrate multiple reward signals.
+For simplicity, we describe the framework using a single augmented sample ($N=2$, easily extendable to multiple samples).
 Let $\tau=\{s_{l}\}_{l\leq L}$ denote the trajectory of the observed positive item $v$, and $\tau^{\prime}=\{s^{\prime}_{l}\}_{l\leq L}$ denote the trajectory of an augmented item $v^{\prime}$.
-Three aspects of reward are defined as:
-  • Augmentation Signal ($r_{a}$): marks whether a sample is observed (positive) or augmented, i.e., $r_{a}=1$ for the ground-truth trajectory $\tau$ and $r^{\prime}_{a}=0$ for the augmented trajectory $\tau^{\prime}$.
-  • Collaborative Score ($r_{c}$): the collaborative score from the same model, i.e., $r_{c}=\hat{y}_{CM}^{v},r^{\prime}_{c}=\hat{y}_{CM}^{v^{\prime}}$.
-  • Token Similarity ($r_{\text{sim}}$): rewards partial correctness by measuring the number of shared tokens with the positive sample.
+Several aspects of reward might be defined as:
+  • Interaction Signal ($r_{a}$): marks the interaction level for items.
+For example, $r_{a}=10$ for the liked item trajectory $\tau$, $r_{a}=1$ for the clicked item trajectory $\tau$, and $r^{\prime}_{a}=0$ for the augmented or unpresented trajectory $\tau^{\prime}$.
+  • Estimated Score ($r_{c}$): the collaborative score from CM, i.e., $r_{c}=\hat{y}_{CM}^{v},r^{\prime}_{c}=\hat{y}_{CM}^{v^{\prime}}$ or the LLM score score (forward probability).
+  • Business Target: customized reward for differnet bussiness targets (Zhou et al., 2025).
+We illustrate with the format reward ($r_{\text{sim}}$), which reveal partial correctness by measuring the number of shared tokens with the positive sample.
 Specifically, for $\tau^{\prime}$, $r^{\prime}_{\text{sim}}=\sum_{l=1}^{L}\mathbb{I}(t_{l}=t^{\prime}_{l})$, where $\mathbb{I}(\cdot)$ is an indicator function.
@@
 For the visualized example, $r^{\prime}_{\text{sim}}=1$, as it has one shared token with the ground truth.
-We sum up these reward signals as the final reward for each trajectory, which can be formally expressed as $\mathcal{R}(\tau)=r_{a}+r_{c}+r_{\text{sim}}$.
-3.3. Training and Inference
-Training.
+3.3. GFN Training Objective
 The training objective of GFlowGR integrates two complementary loss components: the vanilla next-token prediction loss $\mathcal{L}_{\text{GR}}$ for positive samples, the GFlowNet training loss $\mathcal{L}_{\text{GFN}}$ over sampled trajectories.
@@
 (6) $\mathcal{L}_{\text{TB}}(\tau)=\Big(\mathrm{log}\frac{\mathcal{Z}\prod_{l=0}^{L-1}\mathbb{P}_{\text{GR}}(t_{l+1}|U,s_{l})}{\mathcal{R}(s_{L})}\Big)^{2}$
-$\mathcal{F}$, and $\mathcal{Z}$ are learnable GFlowNet components, with $\mathcal{F}(s_{L})=\mathcal{R}(s_{L})$ in Equation (5).
+$\mathcal{F}$ and $\mathcal{Z}$ are learnable GFlowNet components, with $\mathcal{F}(s_{L})=\mathcal{R}(s_{L})$ in Equation (5).
 $\mathcal{L}_{\text{GFN}}$ can be instantiated using either $\mathcal{L}_{\text{DB}}$ or $\mathcal{L}_{\text{TB}}$.
-The complete loss function for $(U,v,\tau)$ is:
-(7) $\mathcal{L}(U,v,\tau)=\mathcal{L}_{\text{GR}}(U,v)+\lambda\sum_{\kappa\in\{\tau,\tau^{\prime}\}}\mathcal{L}_{\text{GFN}}(\kappa)$
-where $\kappa$ represents either the positive trajectory $\tau$ or an augmented trajectory $\tau^{\prime}$, and $\lambda$ is the hyperparameter to balance different objectives.
-To mitigate exposure bias, we weight each trajectory by the inverse disagreement between the augmentation signal $r_{a}$ and the collaborative score $r_{c}$:
-(8) $\mathcal{L}(U,v,\tau)=\mathcal{L}_{\text{GR}}(U,v)+\lambda\sum_{\kappa\in\{\tau,\tau^{\prime}\}}\frac{1}{(r_{a}-r_{c})^{2}}\mathcal{L}_{\text{GFN}}(\kappa)$
-This heuristic weighting strategy downweights examples where the collaborative model exhibits high uncertainty (large $|r_{a}-r_{c}|$), which often correspond to hard negatives or ambiguous interactions (Chen et al., 2023).
+The complete loss function for $(U,v,\{\tau_{n}\}_{n\leq N})$ is:
+(7) $\mathcal{L}(U,v,\{\tau_{n}\}_{n\leq N})=\mathcal{L}_{\text{GR}}(U,v)+\lambda\sum_{n=1}^{N}\mathcal{L}_{\text{GFN}}(\tau_{n})$
+where $\lambda$ is the hyperparameter to balance different objectives.
+3.4. Training and Inference Process
+We delineate the training and inference procedures of GFlowGR in Algorithm 1, with the use of CM.
+Notably, GFlowGR can still work without CM.
+Initially, the employed item tokenizer, generative LLMs for GR frameworks, and the collaborative model for GFlowGR are declared (lines 1 - 3).
+Subsequently, we train the item tokenizer and utilize the trained tokenizer to transform items into identifiers consisting of $L$ tokens (line 5).
+Prior to fine-tuning LLMs with GFlowGR, we declare the essential hyperparameters (line 6).
+For a batch of training samples, we first structure the LLM input (line 8) and convert the item $v$ into the generation trajectory $\tau$ (line 9).
+Next, GFlowGR generates several augmented trajectories (line 10) and computes GFlowNets elements, which include the forward probability, flow, and reward (lines 11 - 13).
+Finally, the LLM parameters are updated by the sum of the SFT loss and the GFlowNets loss (lines 14 - 17).
+During the inference phase, we structure the LLM input, generate the subsequent tokens with the checkpoint, and decode the outcome for recommendation (lines 16 - 18), where GFlowGR is not involved.
+Table 1. The statistics of the preprocessed datasets
+Dataset # Users # Items Sparsity Avg.length
+Yelp 30,431 20,032 99.85% 9.39
+Beauty 22,362 12,101 99.92% 7.87
+Instruments 24,772 9,922 99.99% 7.32
 Table 2.
@@
 For all metrics, the higher is better.
-Model Finetuning Beauty Yelp
-R@5 R@10 N@5 N@10 R@5 R@10 N@5 N@10
-TIGER SFT 0.0354 0.0558 0.0234 0.0299 0.0204 0.0334 0.0132 0.0174
-PPO 0.0020 0.0032 0.0013 0.0017 0.0004 0.0005 0.0003 0.0003
-GRPO 0.0347 0.0543 0.0229 0.0291 0.0193 0.0328 0.0126 0.0169
-DPO 0.0308 0.0483 0.0200 0.0256 0.0100 0.0182 0.0060 0.0087
-S-DPO 0.0314 0.0492 0.0208 0.0265 0.0212 0.0333 0.0137 0.0176
-SPRec 0.0312 0.0487 0.0205 0.0258 0.0103 0.0186 0.0067 0.0090
-IPA 0.0306 0.0481 0.0198 0.0245 0.0101 0.0179 0.0066 0.0093
-GFlowGR-DB 0.0388* 0.0629* 0.0254* 0.0336* 0.0221 0.0345 0.0145 0.0185
-GFlowGR-TB 0.0407* 0.0651* 0.0266* 0.0344* 0.0256* 0.0404* 0.0166* 0.0213*
-LETTER SFT 0.0344 0.0540 0.0227 0.0290 0.0205 0.0341 0.0131 0.0175
-PPO 0.0128 0.0206 0.0087 0.0112 0.0002 0.0004 0.0001 0.0002
-GRPO 0.0384 0.0592 0.0251 0.0317 0.0228 0.0380 0.0146 0.0195
-DPO 0.0358 0.0571 0.0238 0.0307 0.0172 0.0286 0.0115 0.0151
-S-DPO 0.0390 0.0607 0.0259 0.0336 0.0247 0.0389 0.0161 0.0206
-SPRec 0.0370 0.0573 0.0249 0.0315 0.0190 0.0303 0.0124 0.0160
-IPA 0.0367 0.0586 0.0252 0.0320 0.0197 0.0293 0.0129 0.0163
-GFlowGR-DB 0.0407* 0.0630* 0.0266 0.0337 0.0252 0.0399 0.0166 0.0213
-GFlowGR-TB 0.0417* 0.0632* 0.0276* 0.0345* 0.0261* 0.0413* 0.0173* 0.0222*
-Inference.
-The inference process remains unchanged from standard GR frameworks: user prompts are constructed as before, and the fine-tuned LLM generates item identifiers sequentially.
-By design, GFlowGR enhances GR without altering the inference pipeline, ensuring seamless integration into existing systems.
+Model Finetuning Beauty Instruments Yelp
+R@5 R@10 N@5 N@10 R@5 R@10 N@5 N@10 R@5 R@10 N@5 N@10
+TIGER SFT 0.0354 0.0558 0.0234 0.0299 0.0864 0.1071 0.0740 0.0806 0.0204 0.0334 0.0132 0.0174
+GRPO 0.0347 0.0543 0.0229 0.0291 0.0834 0.0988 0.0718 0.0768 0.0193 0.0328 0.0126 0.0169
+DPO 0.0308 0.0483 0.0200 0.0256 0.0803 0.0977 0.0692 0.0753 0.0100 0.0182 0.0060 0.0087
+S-DPO 0.0314 0.0492 0.0208 0.0265 0.0791 0.0930 0.0688 0.0733 0.0212 0.0333 0.0137 0.0176
+SPRec 0.0312 0.0487 0.0205 0.0258 0.0773 0.0903 0.0662 0.0709 0.0103 0.0186 0.0067 0.0090
+IPA 0.0306 0.0481 0.0198 0.0245 0.0769 0.0899 0.0661 0.0707 0.0101 0.0179 0.0066 0.0093
+GFlowGR-DB 0.0388* 0.0629* 0.0254* 0.0336* 0.0868 0.1074 0.0747 0.0807 0.0221 0.0345 0.0145 0.0185
+GFlowGR-TB 0.0407* 0.0651* 0.0266* 0.0344* 0.0882* 0.1093* 0.0747* 0.0815* 0.0256* 0.0404* 0.0166* 0.0213*
+LETTER SFT 0.0344 0.0540 0.0227 0.0290 0.0853 0.1049 0.0734 0.0797 0.0205 0.0341 0.0131 0.0175
+GRPO 0.0384 0.0592 0.0251 0.0317 0.0819 0.0976 0.0709 0.0759 0.0228 0.0380 0.0146 0.0195
+DPO 0.0358 0.0571 0.0238 0.0307 0.0788 0.0960 0.0682 0.0737 0.0172 0.0286 0.0115 0.0151
+S-DPO 0.0390 0.0607 0.0259 0.0336 0.0769 0.0914 0.0674 0.0722 0.0247 0.0389 0.0161 0.0206
+SPRec 0.0370 0.0573 0.0249 0.0315 0.0754 0.0896 0.0655 0.0701 0.0190 0.0303 0.0124 0.0160
+IPA 0.0367 0.0586 0.0252 0.0320 0.0750 0.0891 0.0648 0.0699 0.0197 0.0293 0.0129 0.0163
+GFlowGR-DB 0.0407* 0.0630* 0.0266 0.0337 0.0855 0.1071 0.0738 0.0802 0.0252 0.0399 0.0166 0.0213
+GFlowGR-TB 0.0433* 0.0672* 0.0286* 0.0363* 0.0888* 0.1100* 0.0750* 0.0819* 0.0261* 0.0413* 0.0173* 0.0222*
 4. Experiments
 4.1. Experimental Setups
-Datasets. We used two real-world datasets to evaluate GFlowGR, namely Yelp and Amazon Beauty. The data processing and split are following previous GR works (Wang et al., 2024a; Zheng et al., 2024).
+Datasets.
+In this paper, we conduct comprehensive experiments on three widely-used datasets, i.e., Yelp, Beauty, and Instruments.
+Yelp [^2] https://www.yelp.com/dataset records user check-ins at restaurants, along with corresponding reviews and ratings.
+Beauty and Instruments are subcategories from Amazon Review [^3] https://cseweb.ucsd.edu/~jmcauley/datasets.html#amazon_reviews (McAuley et al., 2015), which include user histories on Amazon products.
+For data preprocessing, we follow the procedures in LETTER (Wang et al., 2024a).
+Items are tokenized based on titles and descriptions for Beauty and Instruments, and based on titles, descriptions, brands, and categories for Yelp.
+For data splitting, we take the last interacted item of each user as the test sample, the second last one as the validation, and iterate the remaining part as training samples, i.e., each interacted data is the target for the preceding user sequence.
+We present the data statistics in Table 1.
 Baselines and Backbones.
-We compare GFlowGR with three categories of LLM fine-tuning baselines: (a) vanilla SFT; (b) on-policy RLFT: PPO (Schulman et al., 2017), GRPO (Shao et al., 2024); (c) off-policy RLFT: DPO (Rafailov et al., 2023), and DPO-based algorithms modified for language-based recommendations, including S-DPO (Chen et al., [n.
+We compare GFlowGR with three categories of LLM fine-tuning baselines: (a) vanilla SFT; (b) on-policy RLFT: GRPO (Shao et al., 2024); (c) off-policy RLFT: DPO (Rafailov et al., 2023), and DPO-based algorithms modified for language-based recommendations, including S-DPO (Chen et al., [n.
 d.]), SPRec (Gao et al., 2025a), IPA (Deng et al., 2025).
@@
 And we install GFlowGR to two backbones: TIGER (Rajput et al., 2023) and LETTER (Wang et al., 2024a).
+Detailed introduction is as follows:
+  • SFT. Supervised fine-tuning, optimizes LLMs with the next token prediction loss as in Equation (1) of the main text.
+  • GRPO (Shao et al., 2024).
+Group Relative Policy Optimization is another on-policy RL algorithm based on the idea of PPO.
+GRPO improves the PPO by updating the model with the relative advantages of a group of actions, reducing the dependency on the absolute reward function.
+  • DPO (Rafailov et al., 2023).
+Direct Policy Optimizaion is an off-policy RL algorithm, which updates the model with off-policy pairwise preference data to simplify training while achieving competitive results in alignment tasks.
+  • S-DPO (Chen et al., [n.
+d.]).
+This is a DPO-like algorithm tailored for language model-based recommendation scenarios.
+Unlike standard DPO, it can utilize multiple negative samples, making it better adapted to the characteristics of recommendation tasks.
+  • SPRec (Gao et al., 2025a).
+This method selects rejected samples for DPO to mitigate item-level and token-level biases during LLM fine-tuning for recommendation tasks.
+Specifically, it updates rejected samples of the off-policy data as negative samples with the highest beam search scores and fine-tunes LLMs via iterative SFT and DPO.
+  • IPA (Deng et al., 2025).
+This method is part of OneRec, which aims to align LLMs with a reward model integrating collaborative knowledge.
+Similar to SPRec, IPA iteratively updates the model via SFT and DPO steps and replaces rejected samples with those having the highest reward at each step.
 Evaluation Metrics.
@@
 To ensure robust results, each test is repeated three times with different random seeds ({42, 43, 44}).
+Implementation Details.
+We conduct all experiments on the same GPU device.
+In terms of the hyperparameter settings, we find the optimal setting for both baselines and GFlowGR.
+The default ensemple of GFlowGR takes the CM-based sampling for trahectory sampler, and the sum of three reward signals as the reward, i.e., $\mathcal{R}(\tau)=r_{a}+r_{c}+r_{\text{sim}}$.
+The training procedure is controlled by Transformer Trainers with early stopping.
+The implementation of baselines is based on TRL (GRPO, DPO) and the released code of papers (S-DPO, SPRec).
+We implement IPA by ourselves, where the reward model is the same collaborative model used for GFlowGR.
+Figure 3. Parameter study on NN and $\lambda$ with TIGER on Beauty. The red dashed line denotes the performance of SFT, which falls outside the plotted range for the figures corresponding to NN.
+Table 3. Ablation study with TIGER on Beauty.
+Model Metrics $r_{a}$ $r_{c}$ $r_{\text{sim}}$ $r_{a}+r_{c}$ $r_{a}+r_{\text{sim}}$ $r_{c}+r_{\text{sim}}$ w/o Ada w/o Traj GFlowGR
+TIGER R@5 0.0403 0.0399 0.0404 0.0399 0.0406 0.0381 0.0386 0.0397 0.0407
+R@10 0.0641 0.0631 0.0634 0.0629 0.0648 0.0600 0.0600 0.0617 0.0651
+N@5 0.0261 0.0263 0.0260 0.0264 0.0260 0.0253 0.0257 0.0261 0.0266
+N@10 0.0341 0.0338 0.0343 0.0337 0.0344 0.0323 0.0326 0.0332 0.0344
+LETTER R@5 0.0387 0.0408 0.0384 0.0398 0.0404 0.0410 0.0424 0.0423 0.0433
+R@10 0.0596 0.0631 0.0580 0.0610 0.0603 0.0622 0.0659 0.0639 0.0672
+N@5 0.0254 0.0265 0.0246 0.0259 0.0263 0.0274 0.0278 0.0274 0.0286
+N@10 0.0322 0.0337 0.0310 0.0327 0.0327 0.0342 0.0354 0.0343 0.0363
 4.2. Main Results
@@
 Comparison with On-policy RLFT.
-On-policy strategies (especially reward-based PPO) fail to achieve competitive results for GR.
-This may stem from out-of-vocabulary item identifiers: LLMs pretrained on general language tasks struggle to explore the OOV item space without a well-defined reward model, leading to poor PPO performance.
-GRPO alleviates this dependency on the reward model via relative advantages.
-(Shao et al., 2024)
+On-policy strategies fail to achieve competitive results for GR.
+This may stem from their sensitivity to the data.
+Although GRPO can also learn from item set with relative advantages.
+(Shao et al., 2024), it maximizes the probability for the positive item, failing to concept the out-of-set items.
+While GFlowGR assigns the probability proportional to the reward without forcing the sum of the item set as 1, possessing the better robustness in term of the augmented item set.
 Comparison with Off-policy RLFT.
@@
 d.]), which is capable of incorporating multiple negative samples, achieves the best results and competes closely with GFlowGR-DB.
-Figure 3. Parameter study on $K$ and $\lambda$ with TIGER on Beauty. The red dashed line denotes the performance of SFT, which falls outside the plotted range for the figures corresponding to $K$.
+This also highlight the urgent need for the ability of learning from the item set rather than single positive item or item pairs.
 4.3. Parameter Analysis
@@
 4.4. Ablation Status
-In this section, we validate the critical components of GFlowGR: the curriculum-style adaptive trajectory sampler, comprehensive reward model, and uncertainty-based trajectory reweighting.
-Table 2 presents ablation results for: all reward signal combinations (six variants, first block of the table, where the name of the variant denotes used reward dignals), removal of adaptive sampling (w/o Ada), omission of Augmented Trajectories (w/o Traj), and removal of reweighting (w/o Wt).
+In this section, we validate the critical components of GFlowGR: the trajectory sampler and the reward model.
+Table 3 presents ablation results for: all reward signal combinations (six variants, first block of the table, where the name of the variant denotes used reward dignals), removal of CM-based adaptive sampling (w/o Ada, using random sampling), omission of Augmented Trajectories (w/o Traj), i.e., $N=1$.
 Key findings include: (a) Loss of any reward signal (augmentation signals, collaborative scores, or token similarities) degrades performance, confirming the necessity of the comprehensive reward model.
 Notably, $r_{a}$ has the most significant impact, while $r_{c}$ contributes marginally due to CM producing less accurate predictions than GR (Rajput et al., 2023; Wang et al., 2024a), leaving to noisy $r_{c}$ signals.
-(b) Removing adaptive sampling or using random sampling reduces effectiveness due to unstable training or low-quality augmented samples.
-(c) Uncertainty-based reweighting contributes to alleviating exposure bias and improving model robustness, validating its role in stabilizing training.
-Table 2. Ablation study with TIGER on Beauty.
-Metrics $r_{a}$ $r_{c}$ $r_{\text{sim}}$ $r_{a}+r_{c}$ $r_{a}+r_{\text{sim}}$ $r_{c}+r_{\text{sim}}$ w/o Ada w/o Traj w/o Wt GFlowGR
-R@5 0.0403 0.0399 0.0404 0.0399 0.0406 0.0381 0.0386 0.0397 0.0388 0.0407
-R@10 0.0641 0.0631 0.0634 0.0629 0.0648 0.0600 0.0600 0.0617 0.0615 0.0651
-N@5 0.0261 0.0263 0.0260 0.0264 0.0260 0.0253 0.0257 0.0261 0.0254 0.0266
-N@10 0.0341 0.0338 0.0343 0.0337 0.0344 0.0323 0.0326 0.0332 0.0327 0.0344
-4.5. In-depth Analysis
-Case Study.
-We visualize a case in Figure 4 to study GFlowGR’s potential in addressing exposure bias.
-The example shows GFlowGR recommending a diverse set of items, including ‘Shampoo’, ‘Scrub’, and ‘Cleanser’ with the ‘Tomatoes’ flavor (bold green), which is a new taste for the user.
-While SFT focuses on ‘Carrots’-flavored ‘Shampoo’ and ‘Conditioner’ from the same brand, which the user recently purchased (bold orange and red).
-This suggests GFlowGR alleviates exposure bias by exploring diverse trajectories, generating exploratory results with different flavors and product.
-Diversity.
-For this experiment, we analyze diversity in top-10 recommendations from GFlowGR and SFT across two dimensions.
-(a) Generation Distribution: We compare models’ item recommendation probabilities against a normal distribution using KL divergence (rather than a uniform distribution because item popularities are inherently different for recommendations (Chen et al., 2023)).
-GFlowGR achieves a closer match (KL=3.13 vs.
-SFT’s 3.24).
-In addition, the generation probability over all items has a 2.2% smaller standard deviation with GFlowGR, indicating a smoother, more diverse distribution.
-(b) Collaborative Scores: We calculate the collaborative score for the recommended items.
-The result shows that GFlowGR’s recommended items have 0.4% lower average collaborative scores but a broader range (0.45 vs.
-SFT’s 0.42) and 3.6% higher standard deviation, showing it prioritizes non-mainstream items while offering greater score diversity.
-These results demonstrate GFlowGR’s enhanced ability to balance relevance and novelty compared to SFT.
-Figure 4. Case study on user-2089 from Beauty with TIGER: We visualized three items from the top-10 list of each method for the brand ‘Yes to’, where the top-1 result of GFlowGR is correct.
+(b) Using random sampling reduces effectiveness due to unstable training or low-quality augmented samples.
+4.5. Deployment
+Since May 2025, GFlowGR has been fully deployed to train our GR models across all Taobao search advertising platforms, including Taobao Mobile, Xianyu, and Taobao Web (as illustrated in Figure 4).
+Offline evaluations, summarized in Table 4, compare GFlowGR against GRPO and SFT on a massive dataset of hundreds of millions of daily requests.
+To maintain confidentiality, training costs are presented relative to the SFT baseline.
+While GRPO’s cost is computationally infeasible, GFlowGR requires approximately double the training time of SFT.
+We note this can be mitigated by scaling resources; doubling the number of GPUs brings GFlowGR’s training time on par with SFT’s.
+In terms of performance, the metrics mirror our findings on public datasets: GFlowGR achieves a significantly larger improvement over the SFT baseline than GRPO does.
+The method’s real-world value was validated in a large-scale online A/B test involving over 100 million user sessions.
+The test revealed statistically significant improvements in core business metrics, including a 0.43% increase in total revenue.
+This lift was particularly pronounced in off-homepage search scenarios, which saw a 1.11% revenue increase.
+These production metrics, collected over 15 consecutive days and verified by Taobao’s internal monitoring systems, confirm the substantial real-world impact.
+Figure 4.
+Deployed business scenarios in Taobao, illustrated with the query “Dress” in the top left red box.
+Advertisements are indicated by blue boxes on the bottom right of each product.
+Due to privacy policy, sensitive information has been obfuscated.
+Method H@20 N@20 Trianing Cost
+[…]
+GRPO 0.427 0.408 -
+[…]
+Table 4. Offline evaluations on production data.
 5. Related Work
 Generative Recommendations.
-Existing GR research can be categorized into three directions: pipeline design, item tokenizer innovations, and LLM decoding enhancement.
-For pipeline design, TIGER (Rajput et al., 2023) pioneered out-of-vocabulary item tokenization and GR workflows, while LC-Rec (Zheng et al., 2024) and ColaRec (Wang et al., 2024b) enhance alignment via tasks or graph collaborative models, with extensions to multi-behavior scenarios (Liu et al., 2024a).
-The second category focuses on plug-in tokenizers (e.g., LETTER (Wang et al., 2024a) for collaborative embeddings) to boost arbitrary GR frameworks.
-The third category optimizes decoding (Lin et al., 2025c; Ding et al., 2024).
-However, LLM fine-tuning for GR remains underexplored, with DPO-based IPA (Deng et al., 2025) as a rare example.
-The proposed GFlowGR is the first to study the exposure bias and diversity.
+According to the composition of GR, existing research can be generally divided into three categories: design of the whole pipeline, innovations on the item tokenizer, and LLM decoding enhancement.
+The representative works for the first category include TIGER (Rajput et al., 2023), LC-Rec (Zheng et al., 2024), and ColaRec (Wang et al., 2024b).
+TIGER (Rajput et al., 2023) is the pioneer in exploring expressing items with out-of-vocabulary tokens and constructs the basic GR pipeline, including the training and inference with item tokenizer and LLMs.
+LC-Rec (Zheng et al., 2024) further enhances the model efficacy by incorporating diverse aligning tasks to align item identifiers to recommendation objectives.
+Similarly, ColaRec (Wang et al., 2024b) introduces the graph-based collaborative models to align the semantics with recommendations.
+Some researchers also explore GR frameworks for other recommendation problems, such as multi-behavior recommendations (Liu et al., 2024a).
+The second category focuses on inventing plug-in item tokenizers to enhance the performance of arbitrary GR frameworks (Zhu et al., 2024; Zheng et al., 2025a; Liu et al., 2024c; Lin et al., 2025b).
+For example, LETTER (Wang et al., 2024a) introduces the collaborative item embeddings and diversity constraints to control the item tokenization process.
+While the last category notices the decoding process for GR and designs the decoding for specific recommendation targets (Lin et al., 2025c; Ding et al., 2024).
+While there are few efforts spent on the LLM fine-tuning specific to GR besides the optimization algorithms for general language-based recommendations (Chen et al., [n.
+d.]; Gao et al., 2025a).
+IPA from OneRec (Deng et al., 2025) is the only LLM fine-tuning work specific to GR.
+The proposed GFlowGR is the pioneer in exploring LLM fine-tuning with the objective of learning from item set.
 Generative Flow Network Applications.
@@
 In this paper, we propose GFlowGR, a GFlowNets-based fine-tuning framework for GR.
-Specifically, we design an adaptive trajectory sampler that leverages collaborative knowledge to select augmented trajectories, enabling the model to explore diverse item spaces.
-We also introduce a comprehensive reward model that integrates multi-faceted signals, including augmentation status, collaborative scores, and token similarities, to provide nuanced feedback for trajectory evaluation.
-During fine-tuning, GFlowGR extends the vanilla SFT loss by incorporating flow-matching objectives, with trajectory-specific losses reweighted based on uncertainty.
-By adaptive sampling, heuristic reweighting, and flow-matching targets, GFlowGR enhances recommendation diversity and addresses the exposure bias challenge inherent in GR frameworks.
+Specifically, we design a trajectory sampler with four strategies to augment the training trajectories, enabling the model to explore diverse item spaces.
+We also introduce a reward model that can integrate multi-faceted signals, exemplified with interaction signal, estimated scores, and bussiness target (illustraed as similarities), to provide nuanced feedback for trajectory evaluation.
+During fine-tuning, GFlowGR extends the vanilla SFT loss by incorporating flow-matching objectives.
+With the ability of learning from the item set, GFlowGR enhances recommendation diversity and quality.
 Extensive experimental results validate the effectiveness of our approach across multiple datasets, backbones, and metrics.